### v2

# Mirror Online Conformal Prediction with Intermittent Feedback

Abstract
Online conformal prediction enables the runtime calibration of a pre-trained artificial intelligence model using feedback on its performance. Calibration is achieved through set predictions that are updated via online rules so as to ensure long-term coverage guarantees. While recent research has demonstrated the benefits of incorporating prior knowledge into the calibration process, this has come at the cost of replacing coverage guarantees with less tangible regret guarantees based on the quantile loss. This work introduces intermittent mirror online conformal prediction (IM-OCP), a novel runtime calibration framework that integrates prior knowledge, while maintaining long-term coverage and achieving sub-linear regret. IM-OCP features closed-form updates with minimal memory complexity, and is designed to operate under potentially intermittent feedback.
Index Terms:
Conformal prediction, calibration, intermittent feedback, online convex optimization
I Introduction
IN safety-critical and high-stakes sequential decision-making processes, such as in robotics [1, 2], wireless communications [3, 4], finance [5], and medicine [6, 7], it is crucial to accurately quantify uncertainty in order to reliably predict the potential outcomes of given actions. Producing reliable uncertainty estimates in these scenarios is often challenging due to the non-stationarity of the data-generating distribution and to the possibility that feedback on the quality of the predictions may be only intermittently available [8].
A widely used approach to quantify uncertainty complements point predictions with prediction sets that satisfy coverage guarantees [9]. Notably, adaptive conformal inference (ACI) [10] converts black-box predictions into prediction sets that enjoy distribution-free, long-term coverage guarantees. This is done by adaptively adjusting prediction set sizes based on feedback about past decisions [11, 12, 13, 14, 15]. Intermittent feedback can be accommodated via intermittent ACI (I-ACI) [16].
In contrast to ACI, the recently introduced Bayesian ACI (B-ACI) adopts a data-centric Bayesian objective that incorporates not only feedback, but also prior information on the prediction scores [17]. When the prior information is well specified, B-ACI can yield prediction sets with lower regret in terms of the quantile loss as compared to ACI. A lower regret with respect to the quantile loss implies better performance in stochastic settings with independent data [17]. However, B-ACI does not guarantee the key calibration requirement of deterministic long-term coverage. Furthermore, it requires storing past data, as well as solving a convex problem at each calibration step.
As summarized in Table I, existing online calibration schemes either cannot incorporate prior information or sacrifice long-term coverage guarantees. To address these limitations, we introduce intermittent mirror online conformal prediction (IM-OCP). IM-OCP builds on online mirror descent (OMD) [18, 19, 20] to integrate prior knowledge, while employing an importance weighting strategy to handle intermittent feedback [21]. This design enables IM-OCP to retain the computational and memory efficiency of ACI, while incorporating prior information. We showcase the merits of IM-OCP for the task of received signal strength indicator (RSSI)-based localization under intermittent feedback [22, 23, 24].
TABLE I: Comparison with the state of the art
Methods
ACI
[10]
I-ACI
[16]
B-ACI
[17]
IB-ACI
IM-OCP
(this work)
long-term coverage
sublinear regret
prior information
memory complexity
closed-form update
intermittent feedback
II Background
II-A Online Calibration via Set Prediction
We consider the problem of online calibration in sequential decision-making problems. At each round $t$, a learner observes an input $X_{t}\in\mathcal{X}$ and aims to predict the corresponding label $Y_{t}\in\mathcal{Y}$. To this end, the predictor assigns a negatively oriented score $s(X_{t},Y)$ to each value $Y\in\mathcal{Y}$, which is assumed to be upper bounded by a constant $B<\infty$, i.e., $s(X,Y)\leq B$. For example, the quadratic loss is defined as $s(X_{t},Y)=|f(X_{t})-Y|^{2}$, where $f(X_{t})$ is the prediction. Using predictor’s output, a prediction set $\mathcal{C}_{t}$ is produced that includes all labels $Y\in\mathcal{Y}$ with a score no larger than a threshold $r_{t}$, i.e.,
$\mathcal{C}_{t}=\mathcal{C}(X_{t},r_{t})=\left\{y\in\mathcal{Y}:s(X_{t},Y)\leq r_{t}\right\}.$ (1)
After producing the decision $\mathcal{C}_{t}$, the system receives feedback in one of the following two forms, from least to most informative: 1) miscoverage error feedback, in which the feedback signal at time $t$ corresponds to the miscoverage error $E_{t}=\mathbbm{1}\{Y_{t}\notin\mathcal{C}_{t}\}$; and 2) score feedback, where the feedback corresponds to the ground truth score $r^{*}_{t}=s(X_{t},Y_{t})$. We assume that feedback is intermittent, so that feedback is available with probability $p_{t}$ at round $t$.
Given a target miscoverage rate $\alpha\in[0,1]$, the goal of calibration is to leverage the outlined feedback information to optimize, in an online fashion, the thresholds $\{r_{t}\}_{t=1}^{T}$ in (1) so as to ensure the long-term coverage guarantee over the time horizon $T$ [10]
$\mathrm{MisCov}(T)=\left|\frac{1}{T}\sum_{t=1}^{T}E_{t}-\alpha\right|\leq AT^{-\gamma}.$ (2)
In (2), the parameters $A$ and $\gamma$ are constants independent of $T$. By (2), the rate of miscoverage errors, $\mathrm{MisCov}(T)$, on an arbitrary sequence $\{(X_{t},Y_{t})\}_{t=1}^{T}$ converges to the target miscoverage level $\alpha$ as the number of steps $T$ tends to infinity.
While the condition (2) applies to an arbitrary sequence $\{(X_{t},Y_{t})\}_{t=1}^{T}$, in some settings one may be justified to assume that the sequence consists of independent and identically distributed (i.i.d.) samples. In this case, the probabilistic coverage condition $\Pr\{Y_{t}\notin\mathcal{C}_{t}\}=\alpha$ can be met for each time $t$ by selecting the threshold $r_{t}$ as the $(1-\alpha)$-quantile of the score distribution. This quantile can be estimated by minimizing the empirical quantile loss
$\ell_{1-\alpha}(r,r^{*}_{t})=(\alpha-\mathbbm{1}\{r<r^{*}_{t}\})(r-r^{*}_{t})$ (3)
as $q_{\alpha}(r^{*}_{1:T})=\min_{u\in\mathbb{R}}\sum_{t=1}^{T}\ell_{1-\alpha}(u,r^{*}_{t})$. In fact, with i.i.d. data, the estimate $q_{\alpha}(r^{*}_{1:T})$ tends to the true $(1-\alpha)$-quantile in the limit $T\rightarrow\infty$.
To assess the capacity of a calibration procedure to perform well also in the case of i.i.d. data, it is thus useful to evaluate the extent to which the thresholds $\{r_{t}\}_{t=1}^{T}$ deviate from the estimated quantile $q_{\alpha}(r^{*}_{1:T})$. Note that the latter can be only evaluated in hindsight, i.e., at time $T$, while the thresholds $r_{t}$ must be produced online at time $t$. This deviation is measured by the regret
$\mathrm{Reg}(T)=\sum_{t=1}^{T}\ell_{1-\alpha}(r_{t},r^{*}_{t})-\min_{u\in\mathbb{R}}\sum_{t=1}^{T}\ell_{1-\alpha}(u,r^{*}_{t}),$ (4)
which amounts to the difference between the cumulative quantile loss of the predicted sequence $\{r_{t}\}_{t=1}^{T}$ and the cumulative loss of the optimal fixed threshold $q_{\alpha}(r^{*}_{1:T})$.
II-B Adaptive Conformal Inference
Assuming that the feedback signal is always available, i.e., $p_{t}=1$ for all $t$, ACI [10] adaptively adjusts the thresholds $r_{t}$ using an online gradient descent (OGD) strategy based on the miscoverage error feedback $E_{t}$. Specifically, the threshold $r_{t}$ is updated using the gradient of the quantile loss (3) as
$r_{t}$ $=r_{t-1}-\eta_{t-1}\nabla\ell_{1-\alpha}(r_{t-1},r_{t-1}^{*})$
$=r_{t-1}-\eta_{t-1}(\alpha-E_{t-1}),$ (5)
where $\eta_{t}>0$ is the step size. Thanks to the regret guarantees of OGD and the restorative nature of the quantile loss [25], given a learning rate sequence $\eta_{t}=c/\sqrt{t}$ with $c>0$ independent of $T$, ACI enjoys both sublinear regret and long-term coverage guarantee (2) with $\gamma=1/2$[14].
ACI has been extended to the intermittent feedback scenario. Denoting as $\text{obs}_{t}$ the binary random variable taking value 1 if feedback $E_{t}$ is observed, I-ACI [16] uses an OGD rule in which the step size is scaled as in
$r_{t}=r_{t-1}-\eta_{t-1}(\alpha-E_{t-1})\frac{\text{obs}_{t-1}}{p_{t-1}}.$ (6)
I-ACI is shown to satisfy expected long-term coverage and to enjoy a sublinear expected regret (see (11) and (14)) [16].
II-C Bayesian Adaptive Conformal Inference
ACI and I-ACI do not allow the integration of prior knowledge about the data distribution in the calibration process. In contrast, B-ACI assumes the availability of a prior distribution $P(r)$, with support $[0,B]$, on the scores $\{r^{*}_{t}\}_{t\geq 1}$. Leveraging this information, along with score feedback, at each time step $t$, the threshold $r_{t}$ is obtained by solving the convex optimization problem
$r_{t}=\arg\min_{r}\left\{\eta_{t}\psi(r)+\frac{1-\eta_{t}}{t-1}\sum_{i=1}^{t-1}\ell_{1-\alpha}(r,r_{i}^{*})\right\},$ (7)
where the convex regularizing function $\psi(r)=\mathbb{E}_{r^{*}\sim P}[\ell_{1-\alpha}(r,r^{*})]$ accounts for prior knowledge.
Unlike ACI and I-ACI, which have constant memory complexity and a simple update rule that can be computed in closed form, B-ACI requires linear memory to store the sequence $\{r^{*}_{t}\}_{t\geq 1}$ of feedback signals, and it requires to solve a convex minimization problem at each round $t$. B-ACI attains sublinear regret, potentially outperforming ACI when the prior knowledge is well specified. However, B-ACI is not guaranteed to satisfy the long-term coverage condition (2).
Similar to ACI, it is possible to extend B-ACI to handle intermittent feedback. This variant, referred to here as IB-ACI, obtains the threshold $r_{t}$ by minimizing the sum of the importance-weighted losses
$r_{t}=\arg\min_{r}\left\{\eta_{t}\psi(r)+\frac{1-\eta_{t}}{t-1}\sum_{i=1}^{t-1}\ell_{1-\alpha}(r,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}\right\}.$ (8)
In Appendix -A, we show that IB-ACI achieves sub-linear expected regret (see (14)).
III Intermittent Mirror Online Conformal Prediction
Given the state of the art summarized in Table I, this paper introduces IM-OCP, a novel prior-dependent calibration scheme with intermittent feedback that exhibits low complexity, sublinear regret, and coverage guarantees. IM-OCP tackles the problem of online calibration via an OMD strategy [18, 19, 20]. This way, IM-OCP retains the simplicity and performance guarantees of first-order methods such as ACI, while allowing for the incorporation of prior information.
For a given prior distribution $P(r)$ defined in $[0,B]$ for the scores $r^{*}$ and a parameter $\sigma>0$, IM-OCP introduce a regularizing function $R(\cdot)$ of the form
$R(r)=\mathbb{E}_{r^{*}\sim P}[\ell_{1-\alpha}(r,r^{*})]+\frac{\sigma}{2}|r|^{2},$ (9)
where the first term follows B-ACI, while the second term ensures the strong convexity of the function $R(\cdot)$. In Appendix -B, we show that the regularizing function (9) is closed, differentiable, $\mu$-strongly convex for some real number $\mu>\sigma$, and $L$-smooth for some real number $L\leq\max_{r\in[0,B]}\{P(r)\}+\sigma$.
Given the regularizing function $R(r)$, the associated Bregman divergence is denoted as $\mathcal{B}_{R}(u,v)=R(u)-R(v)-(\nabla R(v))^{T}(u-v)$. The corresponding mirror map is given by $\mathcal{M}(\cdot)=\nabla R(\cdot)$, with its inverse map defined as $\mathcal{M}^{-1}(\cdot)=(\nabla R)^{-1}(\cdot)$.
At each time $t$, IM-OCP updates the threshold via the OMD rule
$r_{t}=\mathcal{M}^{-1}\left(\mathcal{M}(r_{t-1})-\eta_{t-1}(\alpha-E_{t-1})\frac{\text{obs}_{t-1}}{p_{t-1}}\right).$ (10)
When the feedback $E_{t-1}$ is not observed, and thus we have $\text{obs}_{t-1}=0$, the update (10) corresponds to the identity map. In contract, when the feedback $E_{t-1}$ is observed, the update (10) amounts to a gradient descent step in the dual space associated with the mirror map.
III-A Theoretical Guarantees
We now analyze the performance of IM-OCP, showing that it satisfies expected long-term coverage while also enjoying sublinear expected regret. Proofs are deferred to Supplementary Material.
In the following, IM-OCP coverage is measured via its expected miscoverage error
$\overline{\mathrm{MisCov}}(T)=\left|\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}E_{t}\right]-\alpha\right|,$ (11)
which corresponds to the absolute difference between the expected fraction of miscoverage errors, where the expectation is taken over the feedback observations $\{\text{obs}_{t}\}_{t=1}^{T}$, and the target coverage level $\alpha$.
Similarly to [14], we first establish the following ancillary result stating that the iterates produced by IM-OCP are bounded.
Lemma 1.
For any $r_{1}\in[0,B]$, the iterates produced by IM-OCP satisfy the condition
$r_{t}\in\left[-\frac{\alpha\varpi_{t}}{\mu},B+\frac{(1-\alpha)\varpi_{t}}{\mu}\right]\ \text{for $t>1$},$ (12)
where $\varpi_{t}=\max_{i\in[1,\dots,t-1]}\{\eta_{i}/p_{i}\}$.
Equipped with Lemma 1, we characterize the expected miscoverage error rate of IM-OCP as follows.
Theorem 1.
For any non-increasing step size sequence $\{\eta_{t}\}_{t\geq 1}$, the miscoverage error rate of IM-OCP satisfies the inequality
$\overline{\mathrm{MisCov}}(T)\leq\frac{1}{T\eta_{T}}\left(LB+\frac{L\eta_{1}}{\mu p_{\min}}\right),$ (13)
where $p_{\min}=\min_{t\in[1,2,\cdots,T]}\{p_{t}\}$.
We now prove that IM-OCP enjoys sublinear regret. Similar to (11), we define the IM-OCP regret as
$\overline{\mathrm{Reg}}(T)=\mathbb{E}\left[\sum_{i=1}^{T}{\ell_{\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}}\right]-\min_{u\in\mathbb{R}}\sum_{i=1}^{T}\ell_{\alpha}(u,r_{i}^{*}),$ (14)
Like (4), the quantity in (14) measures the difference between the expected cumulative quantile loss of IM-OCP and the loss of the best threshold chosen in hindsight having observed the entire sequence $\{r_{t}^{*}\}_{t\geq 1}$.
Theorem 2.
For any non-increasing step size sequence $\{\eta_{t}\}_{t\geq 1}$, the regret of IM-OCP in (14) satisfies the inequality
$\overline{\mathrm{Reg}}(T)\leq\frac{D_{T}}{\eta_{T}}+\frac{1}{2\mu p_{\min}}\sum_{i=1}^{T}\eta_{i},$ (15)
where $D_{T}=\max_{t\in\{1,\dots,T\}}\mathcal{B}_{R}(q_{\alpha}(r^{*}_{1:T}),r_{t})$.
Theorems 1 and 2 provide bounds on the expected miscoverage error rate and regret of IM-OCP as a function of the learning rate sequence. The following corollary specializes the miscoverage and regret guarantee under a judicious choice of the learning rate.
Corollary 1.
With a fixed learning rate $\eta=cT^{-\beta}$ or a decaying learning rate $\eta_{t}=ct^{-\beta}$, where $c$ is a positive constant independent of $T$, IM-OCP satisfies the miscoverage guarantee (2) with $\gamma=1-\beta$ and has regret $\mathrm{Reg}(T)=\mathcal{O}(T^{\max\{\beta,1-\beta\}})$.
IV Numerical Results
In this section, we present numerical results for an RSSI-based localization problem assuming intermittent feedback.
IV-A Setting
We consider a multi-building, multi-floor indoor localization problem and adopt the UJIIndoorLoc dataset [26]. The goal is to localize a device positioned in one of three buildings based on the signal power received from $M=520$ access points [22, 23, 24]. Accordingly, the input $X_{t}\in\mathbb{R}^{M}$ corresponds to an $M$-dimensional vector of RSSI values, while the label $Y_{t}=[Y_{t}^{\mathrm{Log}},Y_{t}^{\mathrm{Lat}}]^{T}\in\mathbb{R}^{2}$ represents the longitude and latitude coordinates of the corresponding device to be localized. We apply online calibration to a pre-trained localization model $\hat{Y}_{i}=f_{\text{ELM}}(X_{i})$, implemented using an extreme learning machine with a hidden layer of size 256 [24]. The model is trained using a training set consisting of 1000, 2000, and 8000 samples randomly selected from Buildings 1, 2, and 3, respectively. Since the accuracy of $f_{\text{ELM}}(\cdot)$ is expected to increase in the building index $i$, due to the increasingly larger datasets as $i$ grows, we set a decreasing probability of receiving feedback as $p_{1}=0.5$ for the first building, $p_{2}=0.3$ for the second, and $p_{3}=0.1$ for the third.
IV-B Results
We perform online calibration of the base predictor using I-ACI, B-ACI, and IM-OCP, considering both triangular and truncated Gaussian prior distributions $P(r)$ defined in the interval $[0,B]$. Specifically, for the triangular prior distribution, the mode is set to 0.1, while for the truncated Gaussian prior distribution, the mean is set to 0.1 and the variance to 2. For all schemes, we consider a held-out data sequence of length $T=2400$, with samples randomly selected from each building with equal probability, and set the target miscoverage rate to $\alpha=0.1$.
Figure 1: Miscoverage rate versus the number of test samples ($\alpha=0.1$).
In Fig. 1, we present the miscoverage rate as a function of time index $t$. Both IM-OCP and ACI, which satisfy the miscoverage guarantee (2), exhibit a miscoverage rate that converges to the target value $\alpha$. In contrast, IB-ACI, which does not come with a theoretical coverage guarantee, obtains a final miscoverage rate higher than the target value.
Figure 2: Cumulative quantile loss versus the number of test samples ($\alpha=0.1$).
In Fig. 2, we report the cumulative quantile loss for all algorithms. By leveraging prior information, IM-OCP achieves a lower cumulative loss—and consequently, a smaller regret—compared to I-ACI, while remaining competitive with B-ACI.
Overall, these results demonstrate that IM-OCP satisfies the coverage guarantees while incorporating prior information and achieving a lower regret compared to I-ACI.
V Conclusion
We have proposed IM-OCP, an OCP scheme for intermittent feedback scenarios based on OMD, which uses importance sampling to handle intermittent feedback. IM-OCP allows the incorporation of prior knowledge about the calibration task via the specification of a mirror map, and is proven to achieve both sublinear regret and miscoverage guarantees. The practical effectiveness of IM-OCP is demonstrated through its application to an RSSI-based localization problem. Future work may extend this study to multi-agent online calibration and explore online calibration under robustness requirements to impairments such as missing data.
-A Regret Bound of IB-ACI
We now show that IB-ACI with a decreasing step size $\eta_{t}=t^{-1/2}/2$ enjoys a sublinear expected regret.
We define $h_{t}=\frac{\eta_{t}(t-1)}{1-\eta_{t}}$ and express IB-ACI update rule as follows
$r_{t}=\arg\min_{r}h_{t}\psi(r)+\sum_{i=1}^{t-1}\ell_{1-\alpha}(r,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}.$ (16)
Based on [19, Lemma 2.3], we have
$\sum_{i=1}^{T}\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}-\sum_{i=1}^{T}\ell_{1-\alpha}(u,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ (17)
$\leq h_{T}\psi(u)+\sum_{i=1}^{T}\left(\ell_{1-\alpha}(r_{i},r_{i}^{*})-\ell_{1-\alpha}(r_{i+1},r_{i}^{*})\right)\frac{\text{obs}_{i}}{p_{i}}.$
Noticing that $\ell_{1-\alpha}(r,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ is $\frac{\text{obs}_{i}}{p_{i}}$-Lipschitz, and with the additional assumption that $\psi(r)$ is $\epsilon$-Lipschitz, it follows that $h_{t}\psi(u)$ is $h_{t}\epsilon$-Lipschitz and that
$\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}-\ell_{1-\alpha}(r_{i+1},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ (18)
$\leq\frac{\text{obs}_{i}}{p_{i}}|r_{i}-r_{i+1}|\leq\frac{\text{obs}_{i}}{p_{i}^{2}h_{i}\epsilon}$
By combining (17) and (18), we obtain
$\sum_{i=1}^{T}\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}-\sum_{i=1}^{T}\ell_{1-\alpha}(u,r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}$ (19)
$\leq h_{T}\psi(u)+\sum_{i=1}^{T}\frac{\text{obs}_{i}}{p_{i}^{2}h_{i}\epsilon}$
Taking expectation over the sequence $\{\text{obs}_{i}\}^{T}_{i=1}$
$\mathrm{Reg}(T)\leq\mathbb{E}_{\text{obs}_{i}\sim p_{i}}\left[h_{t}\psi(u)+\sum_{i=1}^{T}\frac{\text{obs}_{i}}{p_{i}^{2}h_{i}\epsilon}\right]$ (20)
$\leq h_{T}\psi(u)+\sum_{i=1}^{T}\frac{1}{p_{i}h_{i}\epsilon}\leq h_{T}\psi(u)+\sum_{i=1}^{T}\frac{1}{p_{\text{min}}h_{i}\epsilon},$
where $p_{\text{min}}=\min\{p_{i}\}$. For the step size choice $\eta=1/(2\sqrt{t})$, we have $h_{t}=\mathcal{O}(\sqrt{t})$ and $1/h_{t}=\mathcal{O}(1/\sqrt{t})$. Thus, IB-ACI achieves sub-linear regret
$\mathrm{Reg}(T)$ $\leq\mathcal{O}\left(\psi(u)\sqrt{T}+\sum_{i=1}^{T}\frac{1}{p_{\text{min}}\epsilon\sqrt{t}})\right)\sim\mathcal{O}(\sqrt{T}).$ (21)
-B Proof of Strongly Convex and Smoothness
The prior regularizer can be expressed as
$R(r)=\mathbb{E}_{r_{i}^{*}\sim P}[\ell_{1-\alpha}(r,r_{i}^{*})]+\frac{\sigma}{2}|r|^{2}$
$=\int_{0}^{B}\ell_{1-\alpha}(r,r^{*})P(r^{*})\text{d}r^{*}+\frac{\sigma}{2}|r|^{2}$ (22)
$=\alpha\int_{0}^{B}(r-r^{*})P(r^{*})\text{d}r^{*}-\int_{r}^{B}(r-r^{*})P(r^{*})\text{d}r^{*}+\frac{\sigma}{2}|r|^{2}.$
where $P(r)$ is the probability density functions of prior distribution $P$ in domain $[0,B]$.
The first derivative of $R(r)$ is
$\nabla R(r)=\int_{0}^{r}P(r^{*})\text{d}r^{*}-(1-\alpha)+\sigma r$ (23)
$\mathop{=}\limits^{(a)}\left\{\begin{array}[]{ll}\mathrm{CDF}_{P}(r)-(1-\alpha)+\sigma r,&0\leq r\leq B,\\ -(1-\alpha)+\sigma r,&r<0,\\ \alpha+\sigma r,&r>B.\end{array}\right.$
where $\mathrm{CDF}_{P}(r)=\int_{0}^{r}P(r^{*})\text{d}r^{*}$ denotes the cumulative distribution function (CDF) of the prior distribution $P(r^{*})$, and (a) follows from the prior distribution $P(r^{*})$ being defined over $[0,B]$, while $P(r^{*})=0$ for $r^{*}\notin[0,B]$, $\mathrm{CDF}_{P}(r^{*})=1$ for $r>B$, and $\mathrm{CDF}_{P}(r^{*})=0$ for $r<0$.
While, the second derivative of $R(r)$ is given by
$\nabla^{2}R(r)=\left\{\begin{array}[]{ll}P(r)+\sigma,&0\leq r\leq B,\\ \sigma,&\text{otherwise}.\end{array}\right.$ (24)
Since $\sigma>0$, we have that $\nabla^{2}R(r)=P(r)+\sigma>\sigma$ for all $r\in[0,B]$. It follows that $R(r)$ is $\mu$-strongly convex with $\mu\geq\sigma$. Similarly, $\nabla^{2}R(r)=P(r)+\sigma\leq\max_{r}\{P(r)\}+\sigma$, which implies that $R(r)$ is $L$-smoothness with $L\leq\max_{r}\{P(r)\}+\sigma$.
Supplementary Material (Appendix)
-C Proof of Lemma 1
Recalling the definition of the mirror map $\mathcal{M}(\cdot)=\nabla R(\cdot)$, the update in the dual space of IM-OCP corresponds to
$\nabla R(r_{t})-\nabla R(r_{t-1})=-\eta_{t-1}(\alpha-E_{t-1})\frac{\text{obs}_{t-1}}{p_{t-1}}.$ (A-1)
Having assumed the $R(r)$ to be $\mu$-strongly convex, it holds
$\left\langle\nabla R(r_{t})-\nabla R(r_{t-1}),r_{t}-r_{t-1}\right\rangle\geq\mu\left\|r_{t}-r_{t-1}\right\|^{2},$ (A-2)
$\left\langle\nabla R(r_{t})-\nabla R(r_{t-1}),r_{t}-r_{t-1}\right\rangle\leq\frac{\left\|\nabla R(r_{t})-\nabla R(r_{t-1})\right\|^{2}}{\mu}$ (A-3)
Leveraging these two inequalities, we now show by contradiction that the value of the threshold $r_{t}$ is upper bounded by $B+(1-\alpha)\varpi_{t}\ {\mu}$ for all $t$. Recall that $\varpi_{t}=\max_{i=1,\dots,t-1}\eta_{i}/p_{i}$ and hence the sequence $\{\varpi_{t}\}_{t\geq 1}$ is non-decreasing. Assume, without loss of generality, that $t$ is the first instant in which the iterate $r_{t}>B+(1-\alpha)\varpi_{t}\ {\mu}$. Given that $\varpi_{t}$ is non-decreasing in $t$, it follows that $r_{t}>r_{t-1}$. From inequality (LABEL:eq:first_ineq_sc) it follows
$r_{t}\leq r_{t-1}-\frac{\eta_{t-1}}{\mu}(\alpha-E_{t-1})\frac{\text{obs}_{t-1}}{p_{t-1}},$ (A-4)
from which we conclude that $E_{t-1}=1$ and $\text{obs}_{t-1}=1$ in order for the inequality $r_{t}>r_{t-1}$ be true. At the same time, given that $E_{t-1}=1$ and $\text{obs}_{t-1}=1$, inequality (LABEL:eq:sec_ineq_sc) implies
$r_{t-1}\geq r_{t}+\frac{\eta_{t-1}(\alpha-1)}{\mu p_{t-1}}>B,$ (A-5)
where the last inequality follows from having initially assumed that $r_{t}>B+(1-\alpha)\varpi_{t}\ {\mu}$. However, this last inequality leads to a contradiction. In fact, since for $r_{t-1}>B$, we have must have $E_{t-1}=0$ and $\text{obs}_{t-1}=1$.
The lower bound follows similarly.
-D Proof of Theorem 1
Note that for IM-OCP it holds that
$\mathbb{E}[\nabla R(r_{t+1})]$ $=\mathbb{E}[\nabla R(r_{\tau})]-\mathbb{E}\left[\sum^{T}_{t=\tau}\eta_{t}(E_{t}-\alpha)\frac{\text{obs}_{t}}{p_{t}}\right]$
$=\mathbb{E}[\nabla R(r_{\tau})]-\sum^{T}_{t=\tau}\eta_{t}(\mathbb{E}[E_{t}]-\alpha).$ (B-1)
Define $\Delta_{1}=1/\eta_{1}$ and $\Delta_{i}=1/\eta_{i}-1/\eta_{i-1}$, and note that $1/\eta_{t}=\sum_{i=1}^{t}\Delta_{i}$. The expected miscoverage rate at time $T$ is given by
$\overline{\mathrm{MisCov}}(T)$ $=\left|\mathbb{E}\left[\frac{1}{T}\sum_{i=1}^{T}\left(\sum_{\tau=1}^{i}\Delta_{\tau}\right)\eta_{i}(E_{i}-\alpha)\right]\right|$ (B-2)
$=\left|\mathbb{E}\left[\frac{1}{T}\sum_{\tau=1}^{T}\Delta_{\tau}\left(\sum_{i={\tau}}^{T}\eta_{i}(E_{i}-\alpha)\right)\right]\right|$
$\leq\frac{\sum_{\tau=1}^{T}|\Delta_{\tau}|}{T}\cdot\left|\mathbb{E}\left[\sum_{i=\tau}^{T}\eta_{i}(E_{i}-\alpha)\right]\right|$
$\leq\frac{\left\|\Delta_{1:T}\right\|_{1}}{T}\cdot\max_{\tau\in\{1,\dots,T\}}\left|\sum_{i=\tau}^{T}\eta_{i}(\mathbb{E}\left[E_{i}\right]-\alpha)\right|$
From (B-1), it holds that
$\sum_{i=\tau}^{T}\eta_{i}(\mathbb{E}\left[E_{i}\right]-\alpha)=\mathbb{E}\left[\nabla R(r_{T+1})-\nabla R(r_{\tau})\right],$ (B-3)
which allows us to rewrite (B-2) as
$\overline{\mathrm{MisCov}}(T)\leq\frac{\left\|\Delta_{1:T}\right\|_{1}}{T}\max_{\tau\in\{1,\dots,T\}}\left|\mathbb{E}\left[\nabla R(r_{T+1})-\nabla R(r_{\tau})\right]\right|.$ (B-4)
Since $R(r)$ is an $L$-smooth function, we obtain
$\max_{\tau\in\{1,\dots,T\}}\left|\mathbb{E}\left[\nabla R(r_{T+1})-\nabla R(r_{\tau})\right]\right|$ (B-5)
$\quad\leq\max_{\tau\in\{1,2,\dots,T\}}L\left|\mathbb{E}\left[r_{T+1}-r_{\tau}\right]\right|LB+\frac{L\eta_{1}}{\mu p_{\min}},$
where the last inequality follows from Lemma 1 and from the assumption of a non-increasing step size sequence $\{\eta_{t}\}_{t\geq 1}$.
By plugging (B-5) into (B-4) and noting that $\left\|\Delta_{1:T}\right\|_{1}=1/\eta_{T}$, we achieve the final result
$\overline{\mathrm{MisCov}}(T)\leq\frac{1}{T\eta_{T}}\left(LB+\frac{L\eta_{1}}{\mu p_{\min}}\right).$ (B-6)
-E Proof of Theorem 2
For any sequence $\{\text{obs}_{t}\}_{t\geq 1}$ and $u\in\mathbb{R}$, online mirror descent satisfies [18, Theorem 6.10]
$\sum_{t=1}^{T}{\ell_{1-\alpha}(r_{t},r_{t}^{*})\frac{\text{obs}_{t}}{p_{t}}}-\sum_{t=1}^{T}{\ell_{1-\alpha}(u,r_{t}^{*})\frac{\text{obs}_{t}}{p_{t}}}$ (C-1)
$\hskip 10.00002pt\leq\frac{D_{T}}{\eta_{T}}+\frac{1}{2\mu}\sum_{t=1}^{T}\eta_{t}\left|\nabla\ell_{1-\alpha}(r_{t},r_{t}^{*})\frac{\text{obs}_{t}}{p_{t}}\right|^{2},$
where $D_{T}=\max_{t\in[1,T]}\mathcal{B}_{R}(u,r_{t})$. Taking expectations over the random variables $\{\text{obs}_{t}\}_{t\geq 1}$ and noting that
$\mathbb{E}\left[\left|\nabla\ell_{1-\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}\right|\right]\leq\mathbb{E}\left[\left|\frac{\text{obs}_{i}}{p_{i}}\right|\right]=1,$ (C-2)
it follows that
$\overline{\mathrm{Reg}}(T)=\mathbb{E}\left[\sum_{i=1}^{T}{\ell_{\alpha}(r_{i},r_{i}^{*})\frac{\text{obs}_{i}}{p_{i}}}\right]-\sum_{i=1}^{T}{\ell_{\alpha}(u,r_{i}^{*})}$ (C-3)
$\leq\frac{D_{T}}{\eta_{T}}+\frac{1}{2\mu}\sum_{i=1}^{T}\frac{\eta_{i}}{p_{i}}\leq\frac{D_{T}}{\eta_{T}}+\frac{1}{2\mu p_{\min}}\sum_{i=1}^{T}\eta_{i},$
where $p_{\min}=\min\{p_{i}\}$. Setting $u=q_{\alpha}(r^{*}_{1:T})$ we obtain the final result.
-F Proof of Corollary 1
For fixed learning rate, by submitting $\eta=cT^{-\beta}$ into miscoverage error rate and regret bound, we have
$\overline{\mathrm{MisCov}}(T)\leq$ $\frac{LB}{c}T^{\beta-1}+\frac{L}{c\mu p_{\min}}T^{-1}\sim\mathcal{O}(AT^{\beta-1}),$ (D-1)
and
$\overline{\mathrm{Reg}}(T)$ $\leq\frac{D_{T}}{c}T^{\beta}+\frac{c}{2\mu p_{\min}}T^{1-\beta}\sim=\mathcal{O}(T^{\max\{\beta,\beta-1\}}).$ (D-2)
Similarly, for decaying learning rate, by submitting $\eta=ct^{-\beta}$ into miscoverage error rate and regret bound, we have
$\overline{\mathrm{MisCov}}(T)\leq\frac{LB\mu{p_{\min}}+Lc}{\mu p_{\min}c}{T^{\beta-1}}\sim\mathcal{O}(AT^{\beta-1}).$ (D-3)
and
$\overline{\mathrm{Reg}}(T)$ $=\frac{D_{T}T^{\beta}}{c}+\frac{c}{2\mu p_{\min}}\sum_{t=1}^{T}t^{-\beta}\sim\mathcal{O}(T^{\max\{\beta,\beta-1\}}).$ (D-4)
By setting $\beta=1/2$, i.e., $\eta_{t}=c\sqrt{t}$, IM-OCP achieves sub-linear regret while ensuring the coverage guarantee.